### v1

# Greed is All You Need: An Evaluation of Tokenizer Inference Methods

Abstract
While subword tokenizers such as BPE and WordPiece are typically used to build vocabularies for NLP models, the method of decoding text into a sequence of tokens from these vocabularies is often left unspecified, or ill-suited to the method in which they were constructed. We provide a controlled analysis of seven tokenizer inference methods across four different algorithms and three vocabulary sizes, performed on a novel intrinsic evaluation suite we curated for English, combining measures rooted in morphology, cognition, and information theory. We show that for the most commonly used tokenizers, greedy inference performs surprisingly well; and that SaGe, a recently-introduced contextually-informed tokenizer, outperforms all others on morphological alignment.
Greed is All You Need: An Evaluation of Tokenizer Inference Methods
[…]
1 Introduction
Modern NLP systems, including large language models (LLMs), typically involve an initial step of mapping raw input text into sequences of subword tokens. These tokens are selected from a large vocabulary of candidates that were produced from algorithms such as Byte-Pair Encoding (BPE; Sennrich et al., 2016), WordPiece Schuster and Nakajima (2012), or UnigramLM Kudo (2018).
This process is referred to as the inference method of tokenization and is critical in nature, as it determines how all words are represented and subsequently modeled. Each inference method offers distinct mappings, and we assert that it is not well-understood how these methods differ in performance. Further, popular implementation packages such as Huggingface Tokenizers, [^1] https://huggingface.co/docs/tokenizers SentencePiece, [^2] https://pypi.org/project/sentencepiece and SubwordNMT [^3] https://github.com/rsennrich/subword-nmt often obfuscate or even restrict the choice of inference methods, making it unclear if the inference-time decoding is suitable with the algorithm used to learn the tokenizer’s vocabulary. Moreover, it is yet to be determined whether such a match is necessary or ideal.
Tokenizerinference ${}_{\text{inference mode}}$ Segmentation
[…]
BPElongest ${}_{\text{longest prefix}}$Ultra modern
[…]
UnigramLMlongest ${}_{\text{longest prefix}}$Un precedent ed
SaGelongest ${}_{\text{longest prefix}}$Inc once iva ble
[…]
Table 1: Examples of words being segmented differently by various tokenizers (vocab size 32,000) using different inference modes on the same vocabulary. Each tokenizer’s default mode is provided on top.
In Table 1 we present examples demonstrating how the prescribed inference methods of BPE, UnigramLM, and SaGe Yehezkel and Pinter (2023) do not necessarily provide the best segmentation for complex English words, even when good segments are available in the vocabulary. BPE’s merge algorithm selects to use the cross-morphemic ‘am’ sequence at an early stage, preventing the consideration of ‘ultra’ and ‘modern’. UnigramLM’s ablative algorithm enabled ‘nprecedented’ (which crosses morpheme boundaries) to remain in its final vocabulary of tokens, while SaGe’s greedy algorithm masks the boundaries of both the prefix ‘In’ and the suffix ‘able’. In all cases, an alternative inference method provides a more morphologically-aligned segmentation over the same vocabulary.
Previous work regarding subword tokenization mostly concerns developing vocabulary construction algorithms Sennrich et al. (2016); Schuster and Nakajima (2012); Kudo (2018); Mielke et al. (2021); Yehezkel and Pinter (2023), finding the optimal vocabulary size (Gowda and May, 2020; Gutierrez-Vasques et al., 2021), building multilingual vocabularies (Liang et al., 2023), and using space positioning in the vocabulary tokens Gow-Smith et al. (2022); Jacobs and Pinter (2022). Others analyzed the effects of vocabularies, finding intricate relations between algorithm and downstream performance Bostrom and Durrett (2020), information theory Zouhar et al. (2023), cognitive plausibility Beinborn and Pinter (2023), or morphological alignment Klein and Tsarfaty (2020); Hofmann et al. (2021, 2022); Gow-Smith et al. (2022).
Research concerning inference methods has been more scarce, and includes examination of random effects on BPE merges Provilkov et al. (2020); Saleva and Lignos (2023) and application of sophisticated search algorithms He et al. (2020). As far as we know, there exists no comprehensive study comparing inference methods across a variety of vocabularies and sizes using diverse metrics.
In this work, we conduct a controlled experiment isolating the effects of inference methods over four tokenizers, introducing an evaluation suite aggregating intrinsic benchmarks from various theoretical realms. [^4] We release our code and data at https://github.com/MeLeLBGU/tokenizers_intrinsic_benchmark. We find that greedy inference methods work surprisingly well for all four vocabularies across morphological and informational metrics. Furthermore, we demonstrate that SaGe yields state-of-the-art performance according to morphological metrics, and that inference methods that minimize token counts perform strongest by cognitive metrics.
2 Inference Methods
Let $\mathcal{V}$ denote a vocabulary of subword tokens and $w$ denote a word (or ‘pretoken’), the output of a pretokenizer. We define $s(\mathcal{V},w):=(t_{1},...,t_{k})$ as a segmentation of $w$ into $k$ subword tokens such that $\forall i,t_{i}\in\mathcal{V}$ and that the concatenation of $t_{1},...,t_{k}$ results in $w$. We use the term segmentation to denote the application of an inference method on a text given a token vocabulary, as well as its result.
Resource Type Size Reference License
LADEC Morphological 7,804 Gagné et al. (2019) CC BY-NC 4.0 DEED
MorphoLex Morphological 12,029 Sánchez-Gutiérrez et al. (2018) CC BY-NC-SA 4.0 DEED
MorphyNet Morphological 219,410 Batsuren et al. (2021) CC BY-SA 3.0 DEED
DagoBert Morphological 279,443 Hofmann et al. (2020) Not specified—citation based
UniMorph Morphological 143,454 Batsuren et al. (2022) CC BY 4.0 DEED
UnBlend Morphological 312 Pinter et al. (2020) GPL-3.0
CompoundPiece Morphological 22,896 Minixhofer et al. (2023) Not specified—citation based
Cognitive data Cognitive 55,867 Beinborn and Pinter (2023) MIT
tokenization-scorer Information Theory Zouhar et al. (2023) Not specified—citation based
Table 2: Size, Reference and License details of the resources in our benchmark.
Current widely-employed tokenization schedules couple together the tokenizer vocabulary with the inference method. However, we advocate for decoupling them, as they are independent processes. Specifically, given a fixed token vocabulary produced from pre-training data, one could subsequently use any applicable inference method for the task at hand. Thus, in our experiments, we use various intrinsic metrics to analyze the impact and performance of the several classes of inference methods:
Greedy inference methods
only consider and produce one token at each step. We define three greedy approaches: Longest prefix, which WordPiece uses by default (Wu et al., 2016), selects the longest token in $\mathcal{V}$ that is a prefix of $w$, and then continues to iteratively segment the remaining text. Longest suffix selects the longest token that is a suffix of $w$ and continues iteratively (Jacobs and Pinter, 2022; Bauwens, 2023). Since this strategy diverges from English Morphology, we consider it an intriguing baseline for assessing the impact of linguistic structure on the inference method. Longest token selects the longest token that is contained in $w$, adds it to the generated segmentation, and then iteratively segments each remaining character sequence. This was proposed by Hofmann et al. (2022) to approximate words by their $k$ longest tokens. They showed that it preserves morphological structure of words and leads to performance gains on some downstream tasks.
Merge rules-based inference methods
begin with a word’s character sequence and iteratively apply token-forming merge rules learnt by the tokenizer at the vocabulary creation phase, until none can be applied. This is BPE’s default inference mode. [^5] While ostensibly also compatible with WordPiece, we found no implementation of the model that provides an ordered list of its merges. In our experiments we test two variants for BPE: The deterministic merge strategy recursively applies the first applicable BPE merge rule by its order in the trained merge list. Dropout (Provilkov et al., 2020) applies each valid merge rule with probability $p$, leading to a regularization effect where rare tokens surface more often and their embeddings can be better trained. It has been shown to improve machine translation performance.
Likelihood-based inference methods
use individual likelihood values assigned to tokens in order to find a segmentation for $w$ where the total likelihood is maximized (Kudo, 2018; He et al., 2020). Default uses likelihood values learned during vocabulary construction and considers the likelihood of a segmentation to be the product of individual likelihoods (from which UnigramLM gets its name). Least tokens assigns a constant likelihood value to all tokens, effectively selecting a segmentation where the number of tokens is minimized.
3 Intrinsic Benchmark
Vocab Inference Morphological Cognitive Rényi Tokens Decoding
method alignment plausibility efficiency per word diff
BPE longest prefix .8584 .3266 .4482 1.4273 .0502
longest suffix .6467 .3170 .4482 1.4286 .0417
longest token .8738 .3302 .4474 1.4261 .0484
least tokens .7544 .3321 .4476 1.4237 .0382
det. merges .6309 .3355 .4482 1.4308
dropout merge .6081 .2925 .4537 1.5793 .1313
WordPiece longest prefix .8488 .3307 .4507 1.4430
longest suffix .6288 .3198 .4502 1.4435 .0656
longest token .8466 .3332 .4500 1.4411 .0216
least tokens .7342 .3306 .4401 1.4319 .0682
UnigramLM longest prefix .9222 .2858 .3400 1.7577 .1187
longest suffix .7520 .2690 .2897 1.7624 .0516
longest token .8845 .2948 .3040 1.7353 .0406
least tokens .8982 .2953 .2969 1.7219 .0328
likelihood .9149 .2937 .2919 1.7314
SaGe longest prefix .9606 .2581 .3217 1.9445
longest suffix .7370 .2471 .2832 1.9615 .1704
longest token .9236 .2671 .3027 1.9236 .0887
least tokens .9125 .2674 .2944 1.8895 .1318
likelihood .9515 .2664 .2937 1.9156 .1168
Table 3: Intrinsic Benchmark results on a vocab size of 40k. ‘Default’ decoding algorithms (used in vocabulary construction) in italics. Not all methods are applicable to all tokenizers. Decoding diff presents the share of pretokens in the MiniPile test set that are differently tokenized using the method, compared with the default.
Some analyses of tokenizers rely on training language models and evaluating their performance on downstream tasks. Using this process to isolate effects of tokenization hyperparameters, such as inference method, is both time- and resource-consuming, as well as unstable due to the introduction of multiple sources of randomness throughout the LM pre-training and fine-tuning phases. Additionally, few measures have been introduced that are intrinsic to vocabularies and their direct application to corpora, and fewer still avoid conflating the measures with the objectives used in the vocabulary construction process itself. As a result, the body of work focused on improving tokenization schemes is still relatively small.
We create and release a benchmark to intrinsically evaluate subword tokenizers. We collected word-level datasets and information measures which have been shown, or hypothesized, to correlate with the performance of language models on various downstream tasks. Details on these resources are provided in Table 2. At present, the benchmark is focused on the English language, although corresponding datasets exist for others as well.
Morphological alignment
[…]
Cognitive Plausibility
We use the benchmark and data from Beinborn and Pinter (2023) to measure the correlation of a tokenizer’s output with the response time and accuracy of human performance on a lexical decision task, predicated on the hypothesis that a good tokenizer struggles with character sequences that humans find difficult, and vice versa. We report the average of the absolute value correlation scores across the four linguistic setups (word/nonword $\times$ accuracy/response time).
Tokens distribution statistics
We report the Rényi efficiency of different segmentations across a corpus Zouhar et al. (2023). This measure penalizes token distributions dominated by either very high- and/or very low-frequency tokens, and was shown to correlate strongly with BLEU scores for machine translation (MT) systems trained on the respective tokenizers. Recent work  (Cognetta et al., 2024) reveals a misalignment between Rényi efficiency and downstream performance in certain cases, reinforcing the necessity of an evaluation suite grounded in diverse domains and disciplines, as advocated in this work. We also measure the average number of tokens per word over a corpus, as a proxy for compression quality. Lastly, we report the proportion of pretokens that are segmented different from the default across this corpus as a reference. We omit the popular measure of character-length distribution of the tokens in the vocabulary, as it does not vary with segmentation strategy.
4 Experiments
We evaluate inference methods for the following tokenizer vocabularies: BPE, UnigramLM, WordPiece and SaGe. We use the train split of the MiniPile (Kaddour, 2023) dataset to construct the tokenizer vocabularies. We train vocabularies of sizes 32,768, 40,960, and 49,152, using the HuggingFace Tokenizers library, with identical pre-tokenization, representing the text at byte level. UnigramLM and SaGe require an initial vocabulary for their top-down algorithms; for the former, we used the default implementation of one million top n-grams, while SaGe was initialized with a 262K-size UnigramLM vocabulary. This initial vocabulary also provided us with token likelihood scores for inference, although a more exact implementation would also incorporate the contextual SaGe objective.
Token distribution statistics measurements were computed over the test split of the MiniPile dataset. We measure the Rényi efficiency using the tokenization-scorer package [^6] https://github.com/zouharvi/tokenization-scorer with $\alpha=2.5$. For each tokenizer, all experiments ran within several minutes on a personal laptop computer.
We present the results on our benchmark for the 40K vocabularies in Table 3. Results for other sizes are presented in Appendix A. A breakdown of individual evaluation subsets is provided in Appendix B.
Inference methods
Within each tokenizer, we find that the default (‘intended’) strategy is often outperformed by others on some measures. We observe a significant difference in morphological alignment when using merge rules-based inference methods. Qualitative analysis showed the findings illustrated in Table 1, where early merge rules such as ‘i-n’, ‘a-m’, or ‘o-n’ cross morphological boundaries. We notice a similar trend for likelihood-based inference, where frequently-used tokens possess very high likelihood values, sometimes exceeding those of the gold-standard segments. We find that the least tokens strategy fares well not only on the token count metric, which is mostly by-design, but also on cognitive measures, suggesting an effect of human preference to minimal word segmentation. Finally, we observe that likelihood-based inference performs poorly in terms of Rényi efficieny, contrary to its stated purpose. Dropout, on the other hand, performs well on this measure, in line with its goal. longest suffix performs poorly across the board, possibly due to the suffixing nature of the English language, which has complementarily been shown to affect character-level sequential modeling (Pinter et al., 2019). Notably, all our key observations are consistent across vocabulary sizes, as shown in Appendix A.
Inter-tokenizer results
Our results align with Bostrom and Durrett (2020)’s finding that BPE is inferior to UnigramLM on morphology alignment. However, we show that some of this gap can be attributed not to the vocabulary but to the inference method. In addition, we find that SaGe is most aligned to morphology by a substantial margin, indicating that its contextualized objective succeeds in retaining meaningful tokens in the vocabulary during ablation. BPE and WordPiece, optimized for compression, unsurprisingly perform well above the likelihood-based vocabularies on the information measures. But, we note that this carries over to the cognitive benchmark as well, supporting Beinborn and Pinter (2023)’s findings.
Finally, we note that the two likelihood-based vocabularies follow the exact same within-vocab trends, and the ones for the two information-based vocabularies are also very close. This highlights the consistency and robustness of our benchmark, although some results are relatively close to each other, which can be expected considering that some inference methods do not change much of the token sequences (see rightmost column of Table 3).
5 Conclusion
In this work, we curated an aggregated benchmark for intrinsic evaluation of subword tokenizers and used it to show the importance of selecting an inference method suited for a vocabulary given a task. Given its computational efficiency, we hope the benchmark can be used in LM training efforts as a fruitful first step to improve tokenization schemes, or to select inference methods on-line. Concretely, our findings suggest that greedy inference is a good choice, especially for morphologically-motivated tasks, even for tokenizers trained on other objectives.
In the future, we plan to examine the correlation between our benchmark and various downstream tasks, as well as expand our experimentation to other languages and new algorithms.
Limitations
Our paper contains evaluation of models in the English language. This was done mostly in order to focus this short paper’s contribution, and to be able to control for as many possibly-confounding variables such as training data. Nevertheless, a more complete followup would have to include attempts to replicate our findings on other languages, aiming for a set as diverse as possible mostly in terms of typology and script.
Our evaluation is limited to intrinsic measures. While this makes development of tokenizers easier, we acknowledge that the body of work correlating success on these measures with performance of downstream models on end-tasks is incomplete.
Ethical Considerations
Details for human annotation for the cognitive benchmark are documented in the source benchmark’s paper Beinborn and Pinter (2023), from which we took the data as-is.
Acknowledgments
We would like to thank Charlie Lovering, Varshini Reddy, and Haoran Zhang for comments on early drafts of this paper. This research was supported in part by the Israel Science Foundation (grant No. 1166/23) and by a Google gift intended for work on Meaningful Subword Text Tokenization.
Appendix A Results on Different Vocabulary Sizes
Table 4 presents benchmark results on 32K-sized and 49K-sized vocabularies.
Vocab Inference Morphological Cognitive Rényi Tokens Decoding
method alignment plausibility efficiency per word diff
BPE-32K longest prefix .8727 .3122 .4600 1.4511 .0581
longest suffix .6496 .3018 .4602 1.4530 .0469
longest token .8883 .3152 .4592 1.4498 .0558
least tokens .7607 .3174 .4595 1.4469 .0426
det. merges .6409 .3201 .4603 1.4551
dropout merge .6149 .2795 .4656 1.6041 .1316
WordPiece-32K longest prefix .7819 .3185 .4630 1.4689
longest suffix .5084 .3089 .4626 1.4698 .0744
longest token .7764 .3212 .4622 1.4667 .0243
least tokens .7394 .3185 .4508 1.4565 .0769
UnigramLM-32K longest prefix .9278 .2855 .3574 1.7803 .1171
longest suffix .7610 .2679 .2961 1.7838 .0516
longest token .8926 .2930 .3103 1.7534 .0395
least tokens .9077 .2937 .3028 1.7418 .0303
likelihood .9206 .2931 .2985 1.7501
SaGe-32K longest prefix .9613 .2610 .3454 1.9502
longest suffix .7449 .2473 .2914 1.9736 .1653
longest token .9348 .2685 .3113 1.9319 .0822
least tokens .9212 .2691 .3035 1.9084 .1247
likelihood .9579 .2679 .3026 1.9246 .1098
BPE-49K longest prefix .8440 .3371 .4391 1.4104 .0444
longest suffix .6438 .3279 .4390 1.4112 .0379
longest token .8637 .3404 .4384 1.4094 .0430
least tokens .7464 .3421 .4385 1.4072 .0351
det. merges .6208 .3461 .4390 1.4137
dropout merge .5967 .2996 .4446 1.5610 .1310
WordPiece-49K longest prefix .7600 .3398 .4413 1.4245
longest suffix .5133 .3309 .4407 1.4247 .0589
longest token .7598 .3421 .4406 1.4228 .0194
least tokens .7261 .3401 .4319 1.4145 .0615
UnigramLM-49K longest prefix .9157 .2818 .3467 1.7432 .1190
longest suffix .7449 .2669 .2849 1.7486 .0516
longest token .8750 .2915 .2994 1.7245 .0416
least tokens .8908 .2926 .2924 1.7098 .0345
likelihood .9095 .2911 .2871 1.7201
SaGe-49K longest prefix .9606 .2566 .3361 1.9414
longest suffix .7355 .2466 .2783 1.9562 .1735
longest token .9200 .2662 .2975 1.9192 .0912
least tokens .9053 .2662 .2893 1.8947 .1353
likelihood .9455 .2651 .2887 $1$1194
Table 4: Aggregated results on 32K and 49K vocabularies.
Appendix B Detailed Results
Table 5 breaks down the results (for 40K) on individual morphological datasets composing our benchmark. Table 6 Provides the same for individual cognitive measures.
Vocab Inference Ladec Morpho- Morphy- Dago- Uni- UnBlend Compound-
Lex Net Bert Morph Piece
BPE longest prefix .9210 .8091 .8511 .8013 .9956 .7404 .8904
longest suffix .9497 .6222 .6524 .7116 .0316 .6095 .9502
longest token .9147 .8125 .8953 .8618 .9705 .7711 .8905
least tokens .9775 .7401 .8303 .8539 .2573 .6489 .9731
det. merges .8160 .6781 .6132 .6195 .3233 .6097 .7568
dropout merge .7666 .6557 .5871 .5953 .3128 .6213 .7178
WordPiece longest prefix .9333 .7625 .9114 .8659 .9963 .5569 .9153
longest suffix .9447 .6005 .6289 .6844 .1059 .4838 .9535
longest token .9275 .7568 .9124 .8765 .9666 .5749 .9112
least tokens .9706 .7132 .8253 .8032 .2670 .5897 .9704
UnigramLM longest prefix .9551 .8800 .9291 .9087 .9973 .8553 .9299
longest suffix .9248 .6387 .8206 .8407 .2777 .8076 .9536
longest token .8855 .7534 .9313 .9378 .9135 .8571 .9130
least tokens .9660 .8015 .9511 .9593 .7218 .9073 .9801
likelihood .9341 .7903 .9645 .9782 .8423 .9205 .9743
SaGe longest prefix .9734 .9422 .9673 .9600 .9973 .9213 .9626
longest suffix .9519 .5996 .7819 .8091 .2403 .8216 .9549
longest token .9420 .8390 .9365 .9418 .9711 .8889 .9457
least tokens .9856 .8394 .9533 .9632 .7269 .9318 .9877
likelihood .9709 .8813 .9809 .9879 .9014 .9492 .9890
Table 5: Results on individual morphological resources.
Vocab Inference Words-RT Words-ACC nonwords-RT nonwords-ACC
BPE longest prefix --.3136 .4035 .4111 --.1784
longest suffix --.3102 .3890 .3987 --.1699
longest token --.3164 .4086 .4130 --.1828
least tokens --.3146 .4083 .4226 --.1828
det. merges --.3285 .4138 .4163 --.1835
dropout merge --.2562 .3505 .3908 --.1726
WordPiece longest prefix --.3198 .4029 .4119 --.1882
longest suffix --.3132 .3863 .4028 --.1770
longest token --.3226 .4067 .4134 --.1902
least tokens --.3146 .4036 .4201 --.1842
UnigramLM longest prefix --.2292 .3391 .3920 --.1827
longest suffix --.2308 .3235 .3645 --.1572
longest token --.2493 .3590 .3904 --.1804
least tokens --.2394 .3582 .3978 --.1860
likelihood --.2424 .3577 .3926 --.1822
SaGe longest prefix --.1924 .2896 .3752 --.1754
longest suffix --.1895 .2801 .3602 --.1585
longest token --.2079 .3047 .3790 --.1767
least tokens --.1978 .3034 .3864 --.1821
likelihood --.2035 .3043 .3797 --.1780
Table 6: A breakdown of cognitive correlation results across vocabularies and inference methods.